OpenStudy (anonymous):

Janelle can get a 24 pack of bottled water for $6.80. HJow much would Janelle have to pay for a 12 pack of bottled water if the ratios are proportional?

OpenStudy (phi):

make a ratio of cost/# of bottles \[ \frac{6.80}{24}\] and another ratio: unknown cost / 12 bottles \( \frac{x}{12}\) and set the ratios equal \[ \frac{6.80}{24} = \frac{x}{12} \] can you solve this ?

OpenStudy (anonymous):

Thank you